Find possible values of a and b that make the statement true. (x) dx + f(x) dx = [* f(x) dx Ja (a, b) = ) x
Answer #1

Property using here is

  • \int_{a}^{b}f\left ( x \right )dx+\int_{b}^{c}f\left ( x \right )dx=\int_{a}^{c}f\left ( x \right )dx

--------------------------------------------------------

\int_{-5}^{1}f\left ( x \right )dx+\int_{1}^{9}f\left ( x \right )dx=\int_{-5}^{9}f\left ( x \right )dx=\int_{a}^{b}f\left ( x \right )dx

\Rightarrow a=-5,b=9

That is

\boldsymbol{\left ( a,b \right )=\left ( -5,9 \right )}

---------------------------------------------------------